# OpenMonoGS-SLAM: Monocular Gaussian Splatting SLAM with Open-set Semantics

### v1

@@ I INTRODUCTION
Simultaneous Localization and Mapping (SLAM) plays a fundamental role in robotics, augmented/virtual reality (AR/VR), and autonomous driving, where understanding spatial layouts and maintaining accurate localization are crucial. In recent years, the focus has expanded beyond purely geometric reconstruction, as spatial AI [SpatialLM, davison2018futuremapping] increasingly emphasizes semantic understanding of 3D environments, an aspect critical for enabling intelligent behavior, meaningful interaction, and effective decision-making.
3D Gaussian Splatting (3DGS) [kerbl3Dgaussians] has emerged as a powerful representation for fast and high-fidelity 3D rendering. It has laid the foundation for several novel SLAM systems that leverage differentiable rasterization for fast and efficient optimization [keetha2024splatam, ha2024rgbd, yan2024gs, peng2024rtg, matsuki2024gaussian, hhuang2024photoslam, wen2025scaffold].
While recent 3DGS-based SLAM methods have shown promising results, many still depend on depth sensors to achieve high-quality reconstruction [keetha2024splatam, ha2024rgbd, yan2024gs, peng2024rtg]. This reliance limits their deployment in scenarios where depth input is unavailable, such as lightweight or low-cost platforms with only RGB cameras. Monocular SLAM offers a more flexible and widely applicable alternative, but existing approaches often struggle to match the geometric accuracy and completeness of their RGB-D counterparts [matsuki2024gaussian, hhuang2024photoslam, wen2025scaffold]. Bridging this performance gap remains a key challenge for making 3DGS-based SLAM practical in real-world, monocular settings.
On the other hand, there is a growing interest in incorporating semantic understanding into 3DGS-based SLAM systems to enable richer scene interpretation beyond geometry. Several recent methods [haghighi2023neural, zhu2024sni, li2024sgs, li2024hier, li2025hier] have explored this direction, aiming to augment 3D representations with semantic labels. However, most of these approaches are limited to closed-set recognition [haghighi2023neural, zhu2024sni, li2024sgs], relying on fixed taxonomies and curated training data, which restricts their ability to handle novel or open-set environments. Furthermore, they often rely on dense depth input [li2024hier, yang2025opengs] for accurate reconstruction and semantic integration, which limits their deployment in monocular or resource-constrained scenarios.
@@
In this work, we present a novel monocular SLAM framework that unifies 3D Gaussian splatting and open-set semantic understanding, named as OpenMonoGS-SLAM. First, building upon the efficient and high-fidelity rendering capabilities of 3DGS, our framework enables accurate and fast 3D reconstruction and camera tracking. Second, relying solely on a monocular RGB input, the proposed method eliminates the need for depth sensors, making it broadly applicable. Third, we move beyond closed-set limitations by integrating open-vocabulary semantic reasoning. To the best of our knowledge, this combination has not been realized in any existing SLAM systems, despite being a direction that aligns closely with the long-term goals of the community.
@@
To achieve our goal, we leverage recent Visual Foundation Models (VFM) [kirillov2023segment, ren2024grounded, radford2021learning, wang2024dust3r, leroy2024grounding] that offer promising zero-shot capabilities and generalization across diverse domains. More specifically, our method begins with a monocular setup that utilizes MASt3R-derived features [leroy2024grounding] for camera tracking and 3D point reconstruction. To obtain semantic cues, we utilize a foundational segmentation model (SAM) [kirillov2023segment] to generate 2D segmentation maps, which are then projected into 3D space. We enforce multi-view consistency to ensure coherent and stable semantic mapping across different viewpoints. Additionally, we employ CLIP [radford2021learning] features to integrate high-level semantic understanding. In summary, our method harnesses the rich generalization capabilities of VFMs to generate semantic features, which are embedded into the 3D Gaussian field for accurate and expressive scene representation.
@@
To further improve the efficiency of OpenMonoGS-SLAM, we propose using a high-dimensional memory bank. We maintain a compact Gaussian feature map throughout the mapping process, and dynamically retrieve relevant high-dimensional CLIP features from the memory bank using an attention mechanism. This enables efficient integration of rich semantic information without inflating the per-Gaussian feature dimensionality. This allows for efficient and expressive representation of semantic information in 3D. Finally, we combine multiple self-supervised learning objectives to effectively learn semantic information without relying on any 3D semantic ground truth, achieving improved performance in both segmentation and mapping.
@@
  • We propose a monocular 3D open-set semantic SLAM framework that integrates VFM-derived semantic features into 3DGS representation. By expressing these features in a multi-scale manner, our approach enables scale-aware and open-set semantic mapping without requiring expensive depth or annotation.
  • We design an efficient and compact semantic fusion strategy that constructs a low-dimensional Gaussian semantic map by embedding 2D features via an attention mechanism over a memory bank of CLIP features. This structure allows efficient semantic retrieval and fusion during mapping, while self-supervised losses enable open-set training without requiring dense annotations.
@@ II-A Visual Foundation Model
Recent advances in visual foundation models have significantly impacted a broad range of computer vision tasks by enabling strong generalization and transferability. These models are typically pretrained on large-scale image datasets using self-supervised or weakly supervised objectives, and are later adapted to a variety of downstream applications. Prominent 2D vision foundation models such as DINO [caron2021dino, oquab2023dinov2] and CroCo [weinzaepfel2022croco, weinzaepfel2023crocov2], learn dense visual representations through contrastive learning or masked image modeling. Trained without explicit human supervision, these models capture rich semantic and geometric information and have been successfully applied to a wide range of tasks, including object discovery [wang2022self_object, zhang2022dino_object], segmentation [kirillov2023segment], dense matching [jiang2024omniglue, edstedt2024roma], and 3D reconstruction [wang2024dust3r, leroy2024grounding, wang2025vggt]. Their dense features serve as general-purpose image descriptors, often used as backbones in more specialized frameworks.
Our method builds upon this line of work by integrating visual representations, promptable segmentation models [kirillov2023segment, radford2021learning] and dense matching frameworks [leroy2024grounding], into a unified system for open-set semantic SLAM. By leveraging the generalization ability of visual foundation models, we aim to minimize task-specific supervision while achieving high performance in complex real-world environments.
@@ II-B Visual SLAM
Dense 3D reconstruction from multi-view images is a long-standing problem in computer vision, fundamental to mapping and scene understanding. The field has progressed from traditional methods, such as Structure-from-Motion [schoenberger2016sfm, pan2024global] and Multi-View Stereo [hartley2003multiple, schoenberger2016mvs], to neural implicit [mildenhall2021nerf] and explicit [kerbl3Dgaussians] approaches, offering improvements in accuracy and representation. However, these methods typically rely on offline processing, limiting their applicability in real-time scenarios.
To address this, recent work has explored integrating neural representations into incremental SLAM pipelines [keetha2024splatam, ha2024rgbd, yan2024gs, peng2024rtg, matsuki2024gaussian, hhuang2024photoslam, wen2025scaffold], enabling online reconstruction while retaining high-quality geometry and appearance. While these approaches have advanced real-time dense mapping, most remain focused on geometry and lack semantic understanding, especially in open-set or unfamiliar environments. Incorporating semantics into SLAM is challenging due to domain shifts, limited annotations, and poor generalization to novel object categories, which restricts the system’s adaptability in real-world scenarios.
Our method addresses these limitations by leveraging visual foundation models for open-set semantic SLAM, unifying promptable semantic segmentation and dense matching representations within a single framework.
Figure 1: Overview of Our Method. Top: Given the previous keyframe and the current frame, MASt3R estimates a point map. We reconstruct a 3D semantic map by augmenting each point with Gaussian attributes and a semantic feature vector. Rendering the 3D Gaussians yields an RGB color and a semantic feature map, which are supervised by the ground truth RGB image and multi-scale masks generated by SAM, respectively. Bottom: When the current frame is selected as a new keyframe, SAM generates instance masks, and masked CLIP features are extracted by applying the masks to the RGB image. These masked CLIP features are used to update the memory bank online and serve as the supervision target for the language-guided loss. The semantic map is further enhanced by memory attention to obtain a high-dimensional semantic map.
@@ III-A1 MASt3R-SLAM
We adopt MASt3R-SLAM [murai2024_mast3rslam] as our backbone model for camera tracking and pointmap estimation. MASt3R employs dense per-pixel features, enabling robust tracking under challenging conditions. We can obtain reliable associations by iteratively optimizing the ray errors. Given the output pointmaps $\mathbf{X}_{i}^{i},\mathbf{X}_{i}^{j}\in\mathbb{R}^{HW\times 3}$, the optimal pixel coordinate in the reference frame $i$, for each target point $\mathbf{x}\in\mathbf{X}_{i}^{j}$ can be found with the following optimization objective,
@@
where $\tilde{\mathbf{X}}_{k,n}^{k},\mathbf{X}_{f,m}^{f}\in\mathbb{R}^{3}$ denote the 3D coordinates of matching features between a frame $f$ and a keyframe $k$, respectively. $\mathbf{T}_{kf}$ represents the relative transformation from frame $f$ to keyframe $k$. $w(\mathbf{q}_{m,n},\sigma_{r}^{2})$ is a confidence-based weighting function where $\mathbf{q}_{m,n}=\sqrt{\mathbf{Q}_{f,m}\mathbf{Q}_{f,n}^{k}}$ is the match confidence score. The Huber norm $\|\cdot\|_{\rho}$ is applied to improve robustness against outliers. Please refer to the original paper for further details [murai2024_mast3rslam].
@@
Similar to [ha2024rgbd], we supervise 3D Gaussians not only with keyframes but also with additional mapping frames, which are used exclusively for optimization. This is necessary since the keyframes selected by MASt3R-SLAM are too sparse to ensure consistent updates across all regions. This representation yields a compact and expressive structure capable of both photorealistic rendering and semantic reasoning.
@@ III-B Multi-Scale Semantic Learning
Recent works have demonstrated the importance of modeling the multi-scale nature of semantic objects for effective 3D scene understanding [kerr2023lerf, kim2024garfield, li2024hier, cen2025segment]. Building on this insight, we aim to guide 3D learning using 2D segmentation outputs from VFMs. To improve generalization to various object sizes and avoid dependence on fixed semantic taxonomies, we adopt a scale-aware strategy. We leverage SAM to produce 2D object masks from each input image. These masks vary in size, reflecting the multi-scale nature of the objects. These 2D masks are then lifted into 3D using depth estimates from MASt3R-SLAM and the known camera intrinsics. We use these semantic information for enhancing multi-view semantic consistency, we encourage the readers to refer to SAGA for more details [cen2025segment].
@@ III-C Language-embedded Semantic Memory Bank
While SAM provides instance-level segmentation, these masks lack semantic meaning without language grounding. [kerr2023lerf] has demonstrated the effectiveness of embedding natural language into radiance fields for semantic understanding. Inspired by this, we produce embeddings for each segmented region using CLIP features to inject language priors into our 3D representation. However, storing raw CLIP embeddings for all Gaussians is prohibitively memory-intensive.
To address this, following the approach used in M3 [zou20253d], we implement a scalable memory bank that stores representative CLIP embeddings. While M3 proposes this mechanism in an offline setting, directly applying it to SLAM is impractical due to the need for online updates. To adapt this idea, we introduce an online mechanism: when a new keyframe is selected, we compute cosine similarity between the current CLIP embedding and the memory bank entries, which store representative CLIP embeddings from previous keyframes. If the similarity is lower than $\tau_{m}$, which means sufficiently different, we append it to the memory bank. This ensures a diverse but compact coverage of the semantic space.
@@ III-D Self-Supervised Learning Objectives
We employ a combination of self-supervised losses to jointly optimize geometry, appearance, and semantics. This synergy between losses allows our model to build expressive 3D maps that are not only photorealistic but also semantically meaningful.
@@
b) Multi-View Semantic Consistency. To enforce semantic consistency across different views, we apply a multi-view contrastive learning objective that pulls semantically similar Gaussians closer and pushes dissimilar ones apart in the feature space. Following SAGA [cen2025segment], the corresponding contrastive loss between two sampled pixel coordinates $\mathbf{p_{1}},\mathbf{p_{2}}$ is written as,
@@
where $s$ denotes the scale index, and $\mathrm{Corr}_{m},\mathrm{Corr}_{f}$ represent mask and feature correspondence, respectively. We sample $|\mathcal{P}|=2,000$ points and use $S=4$ scales. For more details, please refer to SAGA [cen2025segment].
c) Language-Guided Semantic Alignment. To integrate open-set semantic priors from VFMs, we supervise the learned features with CLIP embeddings via a regression objective:
@@
d) Final Loss. We combine all losses into a unified training objective:
@@
By leveraging these complementary loss terms in a self-supervised manner, our model effectively balances geometric accuracy, photometric fidelity, semantic consistency, and language grounding. This enables OpenMono-SLAM to construct robust and interpretable 3D maps from monocular inputs, even in challenging open-set environments.

### v2

@@ I INTRODUCTION
Simultaneous Localization and Mapping (SLAM) plays a fundamental role in robotics, augmented/virtual reality (AR/VR), and autonomous driving, where understanding spatial layouts and maintaining accurate localization are crucial. In recent years, the focus has expanded beyond purely geometric reconstruction, as spatial AI [5] increasingly emphasizes semantic understanding of 3D environments to enable intelligent behavior, meaningful interaction, and effective decision-making. 3D Gaussian Splatting (3DGS) [12] has emerged as a powerful representation for fast and high-fidelity 3D rendering, enabling differentiable rasterization for efficient optimization. This has led to a series of 3DGS-based SLAM systems that achieve real-time tracking and dense reconstruction [11, 8, 34, 19, 9, 33, 25].
Despite these advances, many 3DGS-based SLAM methods still rely on depth sensors to obtain high-quality reconstruction [11, 8, 34]. This reliance limits deployment on lightweight or low-cost platforms where only RGB cameras are available. While monocular alternatives have been explored, they often struggle to match the geometric accuracy and completeness of RGB-D systems [19, 9, 33, 25], leaving a persistent performance gap in practical monocular settings.
Meanwhile, there is growing interest in enriching 3DGS-based SLAM with semantic understanding beyond geometry. Recent methods [18, 16] have explored semantic mapping, yet most remain limited to closed-set recognition [18] with fixed taxonomies and curated data, restricting generalization to novel categories. Moreover, many approaches depend on dense depth input for accurate reconstruction and semantic integration [16, 35], which further constrains applicability in monocular scenarios.
@@
In this work, we present a novel monocular SLAM framework that unifies 3D Gaussian Splatting and open-set semantic understanding, named OpenMonoGS-SLAM. Building on the efficiency of 3DGS, our framework enables accurate and fast 3D reconstruction and camera tracking from monocular RGB input. We further move beyond closed-set limitations by integrating open-vocabulary semantic reasoning. To the best of our knowledge, this combination has not been realized in existing SLAM systems.
@@
To achieve our goal, we leverage recent Visual Foundation Models (VFMs) [14, 23, 15] with strong zero-shot generalization. Specifically, we utilize MASt3R-derived features [15] for camera tracking and 3D point reconstruction, and employ SAM [14] to produce 2D segmentation maps that are lifted into 3D with multi-view consistency for stable semantic mapping. We additionally incorporate CLIP [23] to enable high-level open-vocabulary semantics, embedding these VFM-derived semantic cues into the 3D Gaussian field.
To better handle the inherent multi-scale nature of semantic entities, we convert SAM masks of varying sizes into a scale-aware supervision signal that reduces ambiguity from overlapping masks and encourages semantics that smoothly transition from coarse groupings to fine-grained separation.
@@
To improve efficiency, we maintain a compact Gaussian feature map and retrieve high-dimensional CLIP semantics from a memory bank via attention-based readout, enabling rich semantic fusion without inflating per-Gaussian feature dimensionality. Finally, we combine multiple self-supervised learning objectives to learn semantic information without requiring any 3D semantic ground truth, improving both segmentation and mapping quality.
@@
  • We propose a monocular 3D open-set semantic SLAM framework that integrates VFM-derived semantic features into 3DGS representation. By expressing these features in a multi-scale manner, our approach enables scale-aware and open-set semantic mapping without requiring expensive depth or semantic annotation.
  • We design an efficient semantic fusion strategy that maintains a low-dimensional Gaussian semantic map and uses attention-based readout from a CLIP feature memory bank to retrieve high-dimensional semantics. This structure allows efficient semantic retrieval and fusion during mapping, while self-supervised losses enable open-set training without dense annotations.
  • We show that semantics and geometry mutually reinforce each other in monocular 3DGS SLAM, where semantic consistency improves map stability and geometry accuracy, while better geometry enables reliable semantic lifting.
@@ II-A Visual Foundation Model
Recent advances in visual foundation models have significantly impacted a broad range of computer vision tasks by enabling strong generalization and transferability. These models are typically pretrained on large-scale image datasets using self-supervised or weakly supervised objectives, and are later adapted to a variety of downstream applications. Prominent 2D vision foundation models such as DINO [1] and CroCo [32], learn dense visual representations through contrastive learning or masked image modeling. Trained without explicit human supervision, these models capture rich semantic and geometric information and have been successfully applied to a wide range of tasks, including segmentation [14], dense matching [10, 6], and 3D reconstruction [15, 30]. Their dense features serve as general-purpose image descriptors, often used as backbones in more specialized frameworks.
This line of work suggests that dense correspondence, promptable segmentation, and language-aligned representations can provide strong priors for building semantic 3D understanding with minimal manual supervision. However, directly using VFM outputs in an incremental mapping setting remains challenging, since the signals can be view-dependent, multi-scale, and temporally inconsistent, requiring careful integration across frames and viewpoints.
@@ II-B Visual SLAM
Visual SLAM aims to estimate camera motion while incrementally building a consistent map from streaming images, enabling real-time localization and mapping in robotics and AR/VR [4, 21]. Most classical systems produce sparse or semi-dense maps that are effective for tracking, yet limited for high-fidelity scene modeling and rich scene understanding [21, 7].
In parallel, the mapping component has benefited from advances in dense multi-view reconstruction, ranging from Structure-from-Motion and Multi-View Stereo [26, 27] to neural implicit and explicit representations [20, 12]. While these methods provide detailed geometry and appearance, they often rely on offline optimization, which limits their direct use in real-time SLAM.
To address this, recent work has explored integrating neural representations into incremental SLAM pipelines [11, 8, 34, 19, 9, 33, 25], enabling online reconstruction while retaining high-quality geometry and appearance. Despite this progress, incorporating scalable semantics remains difficult. Many approaches focus primarily on geometry, and semantic extensions are often constrained by depth dependence or fixed label spaces, which limits adaptability to novel categories and open-set environments. Our method addresses these limitations by leveraging visual foundation models for open-set semantic SLAM, unifying promptable semantic segmentation and dense matching representations within a single framework.
Figure 1: Overview of Our Method. Top: Given the previous keyframe and the current frame, MASt3R estimates a point map. The 3D Gaussian scene is initialized by assigning each point with Gaussian attributes and a learnable semantic feature vector. Differentiable rendering produces an RGB image and a semantic feature map, supervised by the ground-truth RGB image and multi-scale SAM masks, respectively. Bottom: When the current frame becomes a new keyframe, SAM predicts instance masks and masked CLIP features are extracted from the RGB image. These features update the memory bank online and provide targets for the language-guided loss. Memory attention then retrieves high-dimensional semantics to refine the Gaussian semantic map.
@@ III-A1 MASt3R-SLAM
We adopt MASt3R-SLAM [22] as our backbone model for camera tracking and pointmap estimation. MASt3R employs dense per-pixel features, enabling robust tracking under challenging conditions. We can obtain reliable correspondences by iteratively optimizing the ray errors. Given the output pointmaps $\mathbf{X}_{i}^{i},\mathbf{X}_{i}^{j}\in\mathbb{R}^{HW\times 3}$, the optimal pixel coordinate in the reference frame $i$, for each target point $\mathbf{x}\in\mathbf{X}_{i}^{j}$ can be found with the following optimization objective,
@@
where $\tilde{\mathbf{X}}_{k,n}^{k},\mathbf{X}_{f,m}^{f}\in\mathbb{R}^{3}$ denote the 3D coordinates of matching features between a frame $f$ and a keyframe $k$, respectively. $\mathbf{T}_{kf}$ represents the relative transformation from frame $f$ to keyframe $k$. $w(\mathbf{q}_{m,n},\sigma_{r}^{2})$ is a confidence-based weighting function where $\mathbf{q}_{m,n}=\sqrt{\mathbf{Q}_{f,m}\mathbf{Q}_{f,n}^{k}}$ is the match confidence score. The Huber norm $\|\cdot\|_{\rho}$ is applied to improve robustness against outliers.
@@
Similar to [8], we supervise 3D Gaussians not only with keyframes but also with additional mapping frames, which are used exclusively for optimization. This is necessary since the keyframes selected by MASt3R-SLAM are too sparse to ensure consistent updates across all regions. This representation yields a compact and expressive structure capable of both photorealistic rendering and semantic reasoning.
@@ III-B Multi-Scale Semantic Learning
Recent works have demonstrated the importance of modeling the multi-scale nature of semantic objects for effective 3D scene understanding [13, 16, 2]. Building on this insight, we aim to guide 3D learning using 2D segmentation outputs from VFMs. To improve generalization to various object sizes and avoid dependence on fixed semantic taxonomies, we adopt a scale-aware strategy. We leverage SAM to produce 2D object masks from each input image. These masks vary in size, reflecting the multi-scale nature of the objects. These 2D masks are then lifted into 3D using depth estimates from MASt3R-SLAM and the known camera intrinsics.
To train scale-aware semantic features, we convert the set of SAM masks into a scale-conditioned supervision signal using $S$ predefined scale levels, inspired by scale-aware contrastive distillation [2]. Concretely, each lifted mask is assigned a 3D scale (e.g., its 3D spatial extent after lifting), and we discretize the scale axis into $S$ levels $\{s\}$ to construct supervision that smoothly transitions from coarse object grouping to fine part-level separation.
For each level $s$, we define a binary pixel identity vector $\mathbf{V}(s,\mathbf{p})\in\{0,1\}^{N}$, where $N$ is the number of lifted masks, indicating which masks are active for pixel $\mathbf{p}$ at that scale. Let $\{\mathbf{M}_{i}\}_{i=1}^{N}$ denote the lifted masks, ordered by their 3D scales from coarse to fine. A key challenge is multi-granularity ambiguity: a pixel may belong to multiple overlapping masks that represent different granularities (e.g., part vs. whole object). To resolve such overlaps in a scale-dependent manner, at scale $s$ we activate mask $\mathbf{M}_{i}$ for pixel $\mathbf{p}$ only if $\mathbf{p}\in\mathbf{M}_{i}$ and there exists no smaller mask $\mathbf{M}_{j}$ such that $\mathbf{p}\in\mathbf{M}_{j}$ and $s\leq\text{size}(\mathbf{M}_{j})<\text{size}(\mathbf{M}_{i})$.
This rule can be interpreted as follows: when $s$ is small (fine levels), many smaller masks satisfy the condition and thus suppress larger overlapping masks, assigning ambiguous regions primarily to fine masks. In contrast, when $s$ is large (coarse levels), fewer smaller masks remain eligible, so larger masks stay active more often, yielding broader, coarse-level groupings. As a result, $\mathbf{V}(s,\mathbf{p})$ provides scale-conditioned membership that reduces ambiguity across granularities.
Finally, we convert the scale-conditioned identity vectors into a binary mask correspondence signal, which is used in our multi-view contrastive objective (Sec. III-D). Specifically, we define the mask correspondence between two pixels $\mathbf{p}_{1}$ and $\mathbf{p}_{2}$ at scale $s$ as
$\mathrm{Corr}_{m}(s,\mathbf{p}_{1},\mathbf{p}_{2})=\mathds{1}\!\left(\mathbf{V}(s,\mathbf{p}_{1})^{\top}\mathbf{V}(s,\mathbf{p}_{2})>0\right),$ (6)
where $\mathds{1}(\cdot)$ denotes the indicator function that returns $1$ if the condition holds and 0 otherwise.
@@ III-C Language-embedded Semantic Memory Bank
While SAM provides instance-level segmentation, these masks lack semantic meaning without language grounding. [13] has demonstrated the effectiveness of embedding natural language into radiance fields for semantic understanding. Inspired by this, we produce embeddings for each segmented region using CLIP features to inject language priors into our 3D representation. However, storing raw CLIP embeddings for all Gaussians is prohibitively memory-intensive.
To address this, similar to M3 [42], we maintain a compact memory bank that stores representative CLIP embeddings to cover the semantic space efficiently. While M3 proposes this mechanism in an offline setting, directly applying it to SLAM is impractical due to the need for online updates. To adapt this idea, we introduce an online mechanism: when a new keyframe is selected, we compute cosine similarity between the current CLIP embedding and the memory bank entries, which store representative CLIP embeddings from previous keyframes. If the similarity is lower than $\tau_{m}$, which means sufficiently different, we append it to the memory bank. This ensures a diverse but compact coverage of the semantic space.
@@ III-D Self-Supervised Learning Objectives
We employ a combination of self-supervised losses to jointly optimize geometry, appearance, and semantics. This synergy between losses allows our model to build expressive 3D maps that are photorealistic and semantically meaningful.
@@
b) Multi-View Semantic Consistency: To enforce semantic consistency across different views, we apply a multi-view contrastive learning objective that pulls semantically similar Gaussians closer and pushes dissimilar ones apart in the feature space. Following SAGA [2], the corresponding contrastive loss between two sampled pixel coordinates $\mathbf{p_{1}},\mathbf{p_{2}}$ is written as,
@@
where $s$ denotes the scale index, and $\mathrm{Corr}_{m},\mathrm{Corr}_{f}$ represent mask and feature correspondence, respectively. We refer to Sec. III-B for the definition of $\mathrm{Corr}_{m}$. We compute $\mathrm{Corr}_{f}$ as the cosine similarity of rendered semantic features:
$\mathrm{Corr}_{f}(s,\mathbf{p}_{1},\mathbf{p}_{2})=\left\langle\mathbf{F}_{\mathbf{p}_{1}},\mathbf{F}_{\mathbf{p}_{2}}\right\rangle.$ (10)
We sample $|\mathcal{P}|=2,000$ points and use $S=4$ scales for multi-scale training. By encouraging consistent semantic features across views, this loss also regularizes the underlying 3D Gaussian map, which empirically leads to more stable camera pose estimation.
c) Language-Guided Semantic Alignment: To integrate open-set semantic priors from VFMs, we supervise the learned features with CLIP embeddings via a regression objective:
@@
d) Final Loss: We combine all losses into a unified training objective:
@@
By leveraging these complementary loss terms in a self-supervised manner, our model effectively balances geometric accuracy, photometric fidelity, semantic consistency, and language grounding. This enables OpenMonoGS-SLAM to construct robust and interpretable 3D maps from monocular inputs, even in challenging open-set environments.